# Moonshot: Optimizing Chain-Based Rotating Leader BFT via Optimistic Proposals

### v1

@@ III-A Protocol Details
We define Simple Moonshot in Figure 1 as a series of event handlers to be run by each node $P_{i}\in\mathcal{V}$. We elaborate on these below.
@@
$P_{i}$ enters view $v$ from some view $v^{\prime}<v$ upon receiving a view $v-1$ block certificate or a view $v-1$ timeout certificate ($P_{i}$ never decreases its local view). Before doing so, it first multicasts this certificate. This ensures that if the first honest node enters $v$ after GST then all honest nodes will enter $v$ or higher within $\Delta$ thereafter, helping our protocol to obtain liveness and reorg resilience. Subsequently, $P_{i}$ updates ${\sf lock}_{i}$ to the highest ranked block certificate that it has received so far, and if ${\sf lock}_{i}$ is not $\mathcal{C}_{v-1}$ then $P_{i}$ unicasts a $\mathsf{status}$ message containing ${\sf lock}_{i}$ to $L_{v}$. We note that $P_{i}$ only updates ${\sf lock}_{i}$ during the view transition process and does not do so after entering the new view, even if it receives a higher ranked block certificate. This ensures that the block certificate reported in a $\mathsf{status}$ message corresponds to its honest sender’s ${\sf lock}_{i}$ for the duration of $v$, meaning that if $L_{v}$ waits to receive status messages from all honest nodes before proposing, then it is guaranteed to extend the block certified by the highest ranked block certificate locked by any honest node. Finally, $P_{i}$ enters $v$, resets ${\sf view\text{-}timer}_{i}$ to $5\Delta$ and starts counting down. If $L_{v}$ is honest and the network is synchronous then $P_{i}$ should enter $v+1$ within $5\Delta$ of entering $v$. If it does not, then it considers the current leader to have failed and so multicasts $\langle\mathsf{timeout},v\rangle_{i}$ to request a view change and prevent the protocol from halting. $P_{i}$ also does this whenever it observes that at least one other honest node has requested a view change for $v$.
@@
Simple Moonshot allows two proposals to be created during view $v$: i) an optimistic proposal for view $v+1$, and; ii) a normal proposal for $v$. In the former case, $L_{v+1}$ multicasts $\langle{\sf opt\text{-}propose},B_{k+1},v+1\rangle$, where $B_{k+1}$ extends $B_{k}$, upon voting for $B_{k}$ in $v$, hoping that $B_{k}$ will become certified. When the protocol is operating in its happy path, $B_{k}$ will indeed become certified, enabling voting for consecutive honest proposals to proceed without delay. In the latter case, $L_{v}$ multicasts $\langle\mathsf{propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),v\rangle$, where $B_{h}$ extends $B_{h-1}$, either upon receiving $\mathcal{C}_{v-1}(B_{h-1})$ within $2\Delta$ time of entering $v$, or after having $\mathcal{C}_{v^{\prime}}(B_{h-1})$ as its highest block certificate after waiting for $2\Delta$ after entering $v$. Since messages are delivered within $\Delta$ time after GST, this $2\Delta$ wait ensures that $L_{v}$ will extend the highest ranked certified block locked by any honest node when it proposes after GST, assisting with liveness and reorg resilience. We require $L_{v}$ to multicast a normal proposal even when it has already multicasted an optimistic proposal to ensure that it always produces a certified block when it proposes after GST. We note that this requirement can be removed from each of our protocols to obtain the corresponding CLSO variant, but doing so naturally sacrifices reorg resilience because the adversary can cause optimistic proposals to fail, even after GST. We discuss how after introducing the remaining protocol rules.
@@
Simple Moonshot has two rules for voting, at most one of which each node may invoke at most once per view. Firstly, $P_{i}$ may $\mathsf{vote}$ for an optimistic proposal containing $B_{k}$ proposed for view $v$ and extending $B_{k-1}$, when locked on $\mathcal{C}_{v-1}(B_{k-1})$. In the best case, $P_{i}$ receives the optimistic proposal containing $B_{k}$ and $\mathcal{C}_{v-1}(B_{k-1})$ simultaneously and so votes for $B_{k}$ immediately upon entering view $v$. Alternatively, if $P_{i}$ receives $\langle\mathsf{propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),v\rangle$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ ranks higher than or equal to ${\sf lock}_{i}$ and $B_{h}$ extends $B_{h-1}$, then it votes for $B_{h}$. Importantly, if $L_{v}$ creates both an optimistic proposal and a normal proposal with the same parent, then since payloads are fixed for a given view, both proposals will contain the same block. This ensures that all honest nodes will vote for the same block, even if they end up using different vote rules.
@@ III-B Analysis
Per Figure 1, Simple Moonshot requires nodes to multicast votes and timeout messages. Each of these messages are $O(1)$ in size, giving these actions a network-level communication complexity of $O(n^{2})$ words per view. Optimistic proposals are likewise $O(1)$ in size, as are normal proposals when threshold signatures are used to compress the vote signatures used to construct block certificates. Accordingly, since the proposal actions are only performed by the leader of the view, they incur $O(n)$ words per view. The forwarding of ${\sf lock}_{i}$ to the next leader upon advancing to a new view incurs a similar cost, assuming threshold signatures, while the multicasting of certificates incurs $O(n^{2})$ communication. Overall then, Simple Moonshot exhibits a network-level communication complexity of $O(n^{2})$ words per view, assuming threshold signatures.
@@ IV Pipelined Moonshot
[t] A Pipelined Moonshot node $P_{i}$ runs the following protocol whilst in view $v$:
$P_{i}$ additionally performs the following actions in any view:
@@
In Pipelined Moonshot, we separate the fallback case of Simple Moonshot’s normal proposal into its own proposal type by enabling $L_{v}$ to create a fallback proposal upon receiving $\mathcal{TC}_{v-1}$. We do so by requiring an honest $P_{i}$ to include its current ${\sf lock}_{i}$ in each $\mathcal{T}$ that it sends. We then allow $L_{v}$ to propose a block that extends the highest ranked $\mathcal{C}$ included in the $\mathcal{T}_{v-1}$ messages used to construct $\mathcal{TC}_{v-1}$.
While this means that $L_{v}$ no longer needs to wait $\Omega(\Delta)$ time before proposing in the fallback path—making Pipelined Moonshot optimistically responsive per Definition 6—it also means that $L_{v}$ may not receive all honest locks before proposing. However, since $\mathcal{TC}$s must be constructed from $12f+$ 1 timeout messages, the highest ranked block certificate in a given $\mathcal{TC}$ necessarily represents the highest lock among at least $f+1$ honest nodes at the time that they sent their timeout messages, which, along with the rules for voting, guarantees that there cannot exist a committable block that was certified for a higher view. This preserves the safety of the protocol. We additionally prevent this modification from breaking liveness by relaxing the voting rule for fallback proposals, allowing $P_{i}$ to vote for the included block if it directly extends the highest ranked block certificate included in the accompanying $\mathcal{TC}$, even if $P_{i}$ is locked on a higher ranked block certificate.
@@ IV-A Protocol Details
As in Simple Moonshot, $P_{i}$ enters view $v$ from some view $v^{\prime}<v$ upon receiving $\mathcal{C}_{v-1}$ or $\mathcal{TC}_{v-1}$. In the former case, as before, it then multicasts $\mathcal{C}_{v-1}$ to assist with reorg resilience and view synchronization. Comparatively, in the latter case $P_{i}$ now unicasts $\mathcal{TC}_{v-1}$ to $L_{v}$ instead of multicasting it. This helps to reduce the communication complexity of the protocol in light of its modified timeout messages, while still ensuring that $L_{v}$ enters $v$ within $\Delta$ of the first honest node doing so after GST. This in turn makes a ${\sf view\text{-}timer}$ of $3\Delta$ sufficient to guarantee the liveness of the protocol, which $P_{i}$ additionally resets regardless of how it enters $v$, and starts counting down. As before, if $P_{i}$ does not advance to a new view before its view timer expires then it multicasts $\langle\mathsf{timeout},v,{\sf lock}_{i}\rangle_{i}$. It likewise multicasts the same message for $v^{\prime\prime}$ upon observing evidence of at least one honest node requesting a view change for $v^{\prime\prime}$ such that $v^{\prime\prime}\geq v$. This latter rule differs from Simple Moonshot and compensates for Pipelined Moonshot’s removal of $\mathcal{TC}$ multicasting.
@@
Pipelined Moonshot consists of three distinct ways to propose a new block in a view; i) an optimistic proposal, ii) a normal proposal, and iii) a fallback proposal. An honest node proposes using at most two of the three methods. The optimistic proposal rule remains the same as in Simple Moonshot and serves the same purpose, allowing voting to proceed without delay when network conditions are favourable. Comparatively, the normal proposal rule now only captures the first case of the same rule in Simple Moonshot: Namely, $L_{v}$ multicasts a normal proposal $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{k-1}),v\rangle$, where $B_{k}$ extends $B_{k-1}$, upon entering view $v$ via $\mathcal{C}_{v-1}(B_{k-1})$. As before, $L_{v}$ does this even if it has already sent an optimistic proposal extending $B_{k-1}$ (which, as before, will necessarily contain $B_{k}$). As in Simple Moonshot, this helps Pipelined Moonshot obtain reorg resilience by ensuring that, after GST, $L_{v}$ will create a proposal that all honest nodes will vote for. Finally, $L_{v}$ multicasts $\langle{\sf fb\text{-}propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),\mathcal{TC}_{v-1},v\rangle$, where $B_{h}$ extends $B_{h-1}$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ is the highest ranked $\mathcal{C}$ in $\mathcal{TC}_{v-1}$, upon entering $v$ via $\mathcal{TC}_{v-1}$. We note that in this case, although we have defined the payload of a block as being fixed for a given view, this restriction can be relaxed in the case of fallback proposals, since the voting rules ensure that if $\mathcal{C}^{f}_{v}(B_{h})$ exists then no other block can be certified for $v$.
@@
In Pipelined Moonshot, $P_{i}$ may vote up to twice in a view; at most once for an optimistic proposal and at most once for either a normal proposal or a fallback proposal. More precisely, $P_{i}$ multicasts $\langle\mathsf{opt\text{-}vote},H(B_{k}),v\rangle_{i}$ for $\langle{\sf opt\text{-}propose},B_{k},v\rangle$, where $B_{k}$ extends $B_{k-1}$, when in view $v$ if it has not yet sent a vote for $v$, or a timeout message for $v-1$ or higher, and has locked $\mathcal{C}_{v-1}(B_{k-1})$. As before, this enables $P_{i}$ to vote for $B_{k}$ immediately upon entering $v$ in the best case. Additionally, $P_{i}$ sends $\langle\mathsf{vote},H(B_{k}),v\rangle_{i}$ for $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{k-1}),v\rangle$ when in $v$ if it has not sent either an $\mathsf{opt\text{-}vote}$ for an equivocating block in view $v$ or a timeout message for view $v$ or higher, and $B_{k}$ extends $B_{k-1}$. Importantly, $P_{i}$ must send this vote if it has already sent an optimistic vote for $B_{k}$. This ensures that $B_{k}$ will be certified when $L_{v}$ is honest and proposes after GST in the case where some honest nodes are unable to send an optimistic vote for $B_{k}$. Otherwise, $P_{i}$ multicasts $\langle\mathsf{fb\text{-}vote},H(B_{h}),v\rangle_{i}$ for $\langle{\sf fb\text{-}propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),\mathcal{TC}_{v-1},v\rangle$ when in view $v$ if it has not sent a timeout message for view $v$ or higher, $B_{h}$ extends $B_{h-1}$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ is the highest ranked block certificate in $\mathcal{TC}_{v-1}$. Notice that this rule allows $P_{i}$ to send a fallback vote for $B_{h}$ after having sent an optimistic vote for an equivocating block, say $B_{k}$. However, since the fallback proposal containing $B_{h}$ can only be valid if it contains $\mathcal{TC}_{v-1}$, at least $f+1$ honest nodes must have sent $\mathcal{T}_{v-1}$ before entering view $v$ and thus will not be able to trigger the optimistic vote rule for $B_{k}$, so $\mathcal{C}^{o}_{v}(B_{k})$ will never exist.

### v2

@@ III-A Protocol Details
We define Simple Moonshot in Figure 1 as a series of event handlers to be run by each node $P_{i}\in\mathcal{V}$. We elaborate below.
@@
$P_{i}$ enters view $v$ from some view $v^{\prime}<v$ upon receiving a view $v-1$ block certificate or a view $v-1$ timeout certificate (i.e. $P_{i}$ never decreases its local view). Before doing so, it first multicasts this certificate. This ensures that if the first honest node enters $v$ after GST then all honest nodes will enter $v$ or higher within $\Delta$ thereafter, helping our protocol to obtain liveness and reorg resilience. Subsequently, $P_{i}$ updates ${\sf lock}_{i}$ to the highest ranked block certificate that it has received so far and if ${\sf lock}_{i}$ is not $\mathcal{C}_{v-1}$ then $P_{i}$ unicasts a $\mathsf{status}$ message containing ${\sf lock}_{i}$ to $L_{v}$. We note that $P_{i}$ only updates ${\sf lock}_{i}$ during the view transition process and does not do so after entering the new view, even if it receives a higher ranked block certificate. This ensures that the block certificate reported in a $\mathsf{status}$ message corresponds to its honest sender’s ${\sf lock}_{i}$ for the duration of $v$, meaning that if $L_{v}$ waits to receive status messages from all honest nodes before proposing, then it is guaranteed to extend the block certified by the highest ranked block certificate locked by any honest node. Finally, $P_{i}$ enters $v$, resets ${\sf view\text{-}timer}_{i}$ to $5\Delta$ and starts counting down. If $L_{v}$ is honest and the network is synchronous then $P_{i}$ should enter $v+1$ within $5\Delta$ of entering $v$. If it does not, then it considers the current leader to have failed and so multicasts $\langle\mathsf{timeout},v\rangle_{i}$ to request a view change and prevent the protocol from halting. $P_{i}$ also does this whenever it observes that at least one other honest node has requested a view change for $v$.
@@
Simple Moonshot allows two proposals to be created during view $v$: i) an optimistic proposal for view $v+1$, and; ii) a normal proposal for $v$. In the former case, $L_{v+1}$ multicasts $\langle{\sf opt\text{-}propose},B_{k+1},v+1\rangle$, where $B_{k+1}$ extends $B_{k}$, upon voting for $B_{k}$ in $v$, hoping that $B_{k}$ will become certified. When the protocol is operating in its happy path after GST, $B_{k}$ will indeed become certified, enabling voting for consecutive honest proposals to proceed without delay. In the latter case, $L_{v}$ multicasts $\langle\mathsf{propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),v\rangle$, where $B_{h}$ extends $B_{h-1}$, either upon receiving $\mathcal{C}_{v-1}(B_{h-1})$ within $2\Delta$ time of entering $v$, or after having $\mathcal{C}_{v^{\prime}}(B_{h-1})$ as its highest block certificate after waiting for $2\Delta$ after entering $v$. Since messages are delivered within $\Delta$ time after GST, this $2\Delta$ wait ensures that $L_{v}$ will extend the highest certified block locked by any honest node when it proposes after GST, assisting with liveness and reorg resilience. We require $L_{v}$ to multicast a normal proposal even when it has already multicasted an optimistic proposal to ensure that it always produces a certified block when it proposes after GST. We note that this requirement can be removed from each of our protocols to obtain the corresponding leader-speaks-once variant, but doing so naturally sacrifices reorg resilience because the adversary can cause optimistic proposals to fail, even after GST. We discuss how after introducing the remaining protocol rules.
@@
Simple Moonshot has two rules for voting, at most one of which each node may invoke at most once per view. Firstly, $P_{i}$ may $\mathsf{vote}$ for an optimistic proposal containing $B_{k}$ proposed for view $v$ and extending $B_{k-1}$, when locked on $\mathcal{C}_{v-1}(B_{k-1})$. In the best case, $P_{i}$ receives the optimistic proposal containing $B_{k}$ and $\mathcal{C}_{v-1}(B_{k-1})$ simultaneously and so votes for $B_{k}$ immediately upon entering view $v$. Alternatively, if $P_{i}$ receives $\langle\mathsf{propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),v\rangle$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ ranks higher than or equal to ${\sf lock}_{i}$ and $B_{h}$ extends $B_{h-1}$, then it votes for $B_{h}$. Importantly, if $L_{v}$ creates both an optimistic proposal and a normal proposal with the same parent, then since payloads are fixed for a given view, both proposals will contain the same block. This ensures that all honest nodes will vote for the same block, even if they use different vote rules.
@@ III-B Analysis
Per Figure 1, Simple Moonshot requires nodes to multicast votes and timeout messages. Each of these messages are $O(1)$ words in size, giving these actions a network-level communication complexity of $O(n^{2})$ words per view. Optimistic proposals are likewise $O(1)$ in size, as are normal proposals when threshold signatures are used to compress the vote signatures used to construct block certificates. Accordingly, since the proposal actions are only performed by the leader of the view, they incur $O(n)$ words per view. The forwarding of ${\sf lock}_{i}$ to the next leader upon advancing to a new view incurs a similar cost, assuming threshold signatures, while the multicasting of certificates incurs $O(n^{2})$ communication. Overall then, Simple Moonshot exhibits a network-level communication complexity of $O(n^{2})$ words per view, assuming threshold signatures.
@@ IV Pipelined Moonshot
A Pipelined Moonshot node $P_{i}$ runs the following protocol whilst in view $v$: 1. Propose. Upon entering $v$ and after executing Advance View and Lock, if $P_{i}$ is $L_{v}$, propose using one of the following rules: (a) Normal Propose. If $L_{v}$ entered $v$ by receiving $\mathcal{C}_{v-1}(B_{k-1})$, multicast $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{k-1}),v\rangle$ such that $B_{k}$ extends $B_{k-1}$. (b) Fallback Propose. If $L_{v}$ entered $v$ by receiving $\mathcal{TC}_{v-1}$, multicast $\langle{\sf fb\text{-}propose},B_{k},\mathcal{C}_{v^{\prime}}(B_{k-1}),\mathcal{TC}_{v-1},v\rangle$ such that $\mathcal{C}_{v^{\prime}}(B_{k-1})$ is ${\sf lock}_{i}$ and $B_{k}$ extends $B_{k-1}$. 2. Vote. $P_{i}$ votes at most twice in view $v$ when the following conditions are met: (a) Optimistic Vote. Upon receiving $\langle{\sf opt\text{-}propose},B_{k},v\rangle$ such that $B_{k}$ extends $B_{k-1}$, if (i) ${\sf timeout\_view}_{i}<v-1$, (ii) ${\sf lock}_{i}=\mathcal{C}_{v-1}(B_{k-1})$ and (iii) $P_{i}$ has not voted in $v$, multicast $\langle\mathsf{opt\text{-}vote},H(B_{k}),v\rangle_{i}$. (b) After executing Advance View and Lock with all embedded certificates, vote once when one of the following conditions are satisfied: i. Normal Vote. Upon receiving $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{h}),v\rangle$, if (i) ${\sf timeout\_view}_{i}<v$, (ii) $B_{k}$ directly extends $B_{h}$ and (iii) $P_{i}$ has not sent an optimistic vote for an equivocating block $B^{\prime}_{k^{\prime}}$ in $v$, multicast $\langle\mathsf{vote},H(B_{k}),v\rangle_{i}$. ii. Fallback Vote. Upon receiving $\langle{\sf fb\text{-}propose},B_{k},\mathcal{C}_{v^{\prime}}(B_{h}),\mathcal{TC}_{v-1},v\rangle$ if (i) ${\sf timeout\_view}_{i}<v$, (ii) $B_{k}$ directly extends $B_{h}$ and (iii) $\mathcal{C}_{v^{\prime}}(B_{h})$ has an equal or greater rank than the highest ranked certificate in $\mathcal{TC}_{v-1}$, multicast $\langle\mathsf{fb\text{-}vote},H(B_{k}),v\rangle_{i}$. 3. Optimistic Propose. Upon voting for $B_{k}$ in view $v$, if $P_{i}$ is $L_{v+1}$, multicast $\langle{\sf opt\text{-}propose},B_{k+1},v+1\rangle$ such that $B_{k+1}$ extends $B_{k}$. 4. Timeout. Upon the expiration of ${\sf view\text{-}timer}_{i}$, if $P_{i}$ has not already sent $\mathcal{T}_{v}$, multicast $\langle\mathsf{timeout},v,{\sf lock}_{i}\rangle_{i}$ and set ${\sf timeout\_view}_{i}=\max({\sf timeout\_view}_{i},v)$. Additionally, upon receiving $f+1$ distinct $\langle\mathsf{timeout},v^{\prime},\_\rangle_{*}$ messages or $\mathcal{TC}_{v^{\prime}}$ such that $v^{\prime}\geq v$ and not having sent $\mathcal{T}_{v^{\prime}}$, multicast $\langle\mathsf{timeout},v^{\prime},{\sf lock}_{i}\rangle_{i}$ and set ${\sf timeout\_view}_{i}=\max({\sf timeout\_view}_{i},v^{\prime})$. 5. Advance View. $P_{i}$ enters $v^{\prime}$ where $v^{\prime}>v$ using one of the following rules: - Upon receiving $\mathcal{C}_{v^{\prime}-1}(B_{h})$. Also, multicast $\mathcal{C}_{v^{\prime}-1}(B_{h})$. - Upon receiving $\mathcal{TC}_{v^{\prime}-1}$. Also, unicast $\mathcal{TC}_{v^{\prime}-1}$ to $L_{v^{\prime}}$. Finally, reset ${\sf view\text{-}timer}_{i}$ to $3\Delta$ and start counting down. $P_{i}$ additionally performs the following actions in any view: 1. Lock. Upon receiving $\mathcal{C}_{v}(B_{k})$ in any protocol message whilst having ${\sf lock}_{i}=\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ such that $v>v^{\prime}$, set ${\sf lock}_{i}$ to $\mathcal{C}_{v}(B_{k})$. 2. Direct Commit. Upon receiving $\mathcal{C}_{v-1}(B_{k-1})$ and $\mathcal{C}_{v}(B_{k})$ such that $B_{k}$ extends $B_{k-1}$, commit $B_{k-1}$. 3. Indirect Commit. Upon directly committing $B_{k-1}$, commit all of its uncommitted ancestors.
@@
In Pipelined Moonshot, we separate the fallback case of Simple Moonshot’s normal proposal into its own proposal type by enabling $L_{v}$ to create a fallback proposal extending its ${\sf lock}$ upon entering $v$ via $\mathcal{TC}_{v-1}$. While this means that $L_{v}$ no longer needs to wait $\Omega(\Delta)$ time before proposing in the fallback path—making Pipelined Moonshot optimistically responsive per Definition 6—it also means that $L_{v}$ may not receive the locks of all honest nodes before proposing. However, since $\mathcal{TC}$s must be constructed from $12f+$ 1 timeout messages, which in turn must now include the sender’s ${\sf lock}$, $L_{v}$ must process the locks of at least $f+1$ honest nodes before creating its proposal. Consequently, $L_{v}$’s ${\sf lock}$ is guaranteed to have a rank at least as great as the highest the highest ranked ${\sf lock}$ among these nodes at the time that they sent their timeout messages. This, along with the rules for voting, guarantees that there cannot exist a committable block with a higher height than $L_{v}$’s ${\sf lock}$. This helps to preserve safety in light of the additional modification that we make to preserve liveness, which we do by allowing $P_{i}$ to vote for $\langle{\sf fb\text{-}propose},B_{k},\mathcal{C}_{v^{\prime}}(B_{h}),\mathcal{TC}_{v-1},v\rangle$ even if it has ${\sf lock}_{i}>\mathcal{C}_{v^{\prime}}(B_{h})$, given $B_{k}$ directly extends $B_{h}$ and $\mathcal{C}_{v^{\prime}}(B_{h})$ has a rank at least as great as the highest ranked block certificate included in $\mathcal{TC}_{v-1}$.
@@ IV-A Protocol Details
As in Simple Moonshot, $P_{i}$ enters view $v$ from some view $v^{\prime}<v$ upon receiving $\mathcal{C}_{v-1}$ or $\mathcal{TC}_{v-1}$. In the former case, as before, it then multicasts $\mathcal{C}_{v-1}$ to assist with reorg resilience and view synchronization. Comparatively, in the latter case $P_{i}$ now unicasts $\mathcal{TC}_{v-1}$ to $L_{v}$ instead of multicasting it. This helps to reduce the communication complexity of the protocol in light of its modified timeout messages, while still ensuring that $L_{v}$ enters $v$ within $\Delta$ of the first honest node doing so after GST. This in turn makes a ${\sf view\text{-}timer}$ of $3\Delta$ sufficient to guarantee the liveness of the protocol (which can be further optimized under crashed leaders, as explained in Appendix D), which $P_{i}$ additionally resets regardless of how it enters $v$, and starts counting down. As before, if $P_{i}$ does not advance to a new view before its view timer expires then it multicasts $\langle\mathsf{timeout},v,{\sf lock}_{i}\rangle_{i}$. It likewise multicasts the same message for $v^{\prime\prime}$ upon observing evidence of at least one honest node requesting a view change for $v^{\prime\prime}$ such that $v^{\prime\prime}\geq v$. This latter rule differs from Simple Moonshot and compensates for Pipelined Moonshot’s removal of $\mathcal{TC}$ multicasting.
@@
Pipelined Moonshot consists of three distinct ways to propose a new block in a view; i) an optimistic proposal, ii) a normal proposal, and iii) a fallback proposal. An honest node proposes using at most two of the three methods. The optimistic proposal rule remains the same as in Simple Moonshot and serves the same purpose, allowing voting to proceed without delay when network conditions are favourable. Comparatively, the normal proposal rule now only captures the first case of the same rule in Simple Moonshot: Namely, $L_{v}$ multicasts a normal proposal $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{k-1}),v\rangle$, where $B_{k}$ extends $B_{k-1}$, upon entering view $v$ via $\mathcal{C}_{v-1}(B_{k-1})$. As before, $L_{v}$ does this even if it has already sent an optimistic proposal extending $B_{k-1}$ (which, as before, will necessarily contain $B_{k}$). As in Simple Moonshot, this helps Pipelined Moonshot obtain reorg resilience by ensuring that, after GST, $L_{v}$ will create a proposal that all honest nodes will vote for. Finally, $L_{v}$ multicasts $\langle{\sf fb\text{-}propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),\mathcal{TC}_{v-1},v\rangle$, where $B_{h}$ extends $B_{h-1}$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ is ${\sf lock}_{i}$, upon entering $v$ via $\mathcal{TC}_{v-1}$. Importantly, since $L_{v}$ only attempts this proposal after executing the Lock rule, $\mathcal{C}_{v^{\prime}}(B_{h-1})$ is guaranteed to have a rank greater than or equal to that of the highest ranked certificate included in $\mathcal{TC}_{v-1}$. We note that in the case of fallback proposals, the requirement that blocks created for the same view must contain the same payload can be relaxed because the voting rules ensure that if $\mathcal{C}^{f}_{v}(B_{h})$ exists then no other block can be certified for $v$.
@@
In Pipelined Moonshot, $P_{i}$ may vote up to twice in a view; at most once for an optimistic proposal and at most once for either a normal proposal or a fallback proposal. More precisely, $P_{i}$ multicasts $\langle\mathsf{opt\text{-}vote},H(B_{k}),v\rangle_{i}$ for $\langle{\sf opt\text{-}propose},B_{k},v\rangle$, where $B_{k}$ extends $B_{k-1}$, when in view $v$ if it has not yet sent a vote for $v$, or a timeout message for $v-1$ or higher, and has locked $\mathcal{C}_{v-1}(B_{k-1})$. As before, this enables $P_{i}$ to vote for $B_{k}$ immediately upon entering $v$ in the best case. Additionally, $P_{i}$ sends $\langle\mathsf{vote},H(B_{k}),v\rangle_{i}$ for $\langle\mathsf{propose},B_{k},\mathcal{C}_{v-1}(B_{k-1}),v\rangle$ when in $v$ if it has not sent either an $\mathsf{opt\text{-}vote}$ for an equivocating block in view $v$ or a timeout message for view $v$ or higher, and $B_{k}$ extends $B_{k-1}$. Importantly, $P_{i}$ must send this vote if it has already sent an optimistic vote for $B_{k}$. This ensures that $B_{k}$ will be certified when $L_{v}$ is honest and proposes after GST in the case where some honest nodes are unable to send an optimistic vote for $B_{k}$. Otherwise, $P_{i}$ multicasts $\langle\mathsf{fb\text{-}vote},H(B_{h}),v\rangle_{i}$ for $\langle{\sf fb\text{-}propose},B_{h},\mathcal{C}_{v^{\prime}}(B_{h-1}),\mathcal{TC}_{v-1},v\rangle$ when in view $v$ if it has not sent a timeout message for view $v$ or higher, $B_{h}$ extends $B_{h-1}$ and $\mathcal{C}_{v^{\prime}}(B_{h-1})$ has a rank greater than or equal to that of the highest ranked block certificate in $\mathcal{TC}_{v-1}$. Notice that this rule allows $P_{i}$ to send a fallback vote for $B_{h}$ after having sent an optimistic vote for an equivocating block, say $B_{k}$. However, since the fallback proposal containing $B_{h}$ can only be valid if it contains $\mathcal{TC}_{v-1}$, at least $f+1$ honest nodes must have sent $\mathcal{T}_{v-1}$ before entering view $v$ and thus will not be able to trigger the optimistic vote rule for $B_{k}$, so $\mathcal{C}^{o}_{v}(B_{k})$ will never exist.